HIGH RESOLUTION PHOTOELECTRON SPECTROSCOPY OF Au_{2}^{−} and Au_{4}^{−} BY PHOTOELECTRON IMAGING

\underline{IKER LEON}, ZHENG YANG, and LAI-SHENG WANG, DEPARTMENT OF CHEMISTRY, BROWN UNIVERSITY, PROVIDENCE, RI 02912, USA.

We report high resolution photoelectron spectra of Au_{2}^{−} and Au_{4}^{−} obtained with a newly-built photoelectron imaging apparatus. Gold anions are produced by laser vaporization and the desired specie is mass selected and focused into the collinear velocity-map imaging (VMI) lens assembly. The design of the imaging lens has allowed us to obtain less than 0.9% energy resolution for high kinetic energy electrons ( > 1eV) while maintaining wavenumber resolution for low kinetic energy electrons. Although gold dimer and tetramer have been studied in the past, we present spectroscopic results under high resolution. For Au_{2}^{−}, we report high resolution spectra with an accurate determination of the electron affinity together with a complete vibrational assignment, for both the anion and neutral ground states, while for Au_{4}^{−}, we are able to resolve a low frequency mode and obtain accurately the adiabatic detachment energy.